# Gaitsgory’s central functor and the Arkhipov–Bezrukavnikov equivalence in mixed characteristic

### v3

@@ 3.3. Cohomology of Wakimoto filtered objects
By definition ${\mathcal{F}}\ast{\mathscr{I}}_{{\bar{\nu}}}({\Lambda})\cong R{\operatorname{pr}}_{2,\ast}({\operatorname{pr}}_{1}^{\ast}({\mathcal{F}})\otimes^{L}_{\Lambda}\pi^{\ast}\nabla_{\bar{\nu}}({\Lambda}))$ and thus we get
@@
This result is analogous to [AR, Proposition 4.5.4]. The existence of a canonical isomorphism follows from Corollary 3.21 by using a filtration of $\mathcal{F}$ by Wakimoto sheaves. Note that the associated graded of ${\mathcal{F}}$ is perverse, so we conclude that, whenever $\bar{\nu}_{1}\preceq\bar{\nu}_{2}$, then the cohomology complexes $R\Gamma({\mathrm{Fl}}_{\mathcal{I}},\mathrm{gr}_{\bar{\nu}_{i}}({\mathcal{F}}))$ sit in different degrees with the same parity. This implies that the connecting homomorphisms of the associated long exact sequences vanish. ∎
@@
By Braden’s theorem, the left side of (3.50) naturally identifies with cohomology supported at the corresponding $LU$-orbit ${\mathcal{S}}_{t_{\bar{\nu}}w}$. Since ${\mathcal{S}}_{t_{\bar{\nu}}}$ contains ${\mathrm{Fl}}_{{\mathcal{I}},t({\lambda})}$ by Lemma 2.12, the pullback of ${\mathcal{S}}_{t_{\bar{\nu}}w}$ to ${\mathrm{Fl}}_{{\mathcal{I}},\leq t({\lambda})}\tilde{\times}{\mathrm{Fl}}_{\mathcal{I}}$ identifies with $L^{+}{\mathcal{U}}\bar{\nu}(\pi)L^{+}{\mathcal{U}}\times^{L^{+}{\mathcal{U}}}{\mathcal{S}}_{w}$. By abuse of notation, we denote the latter scheme by ${\mathrm{Fl}}_{t({\lambda})}\tilde{\times}{\mathcal{S}}_{w}$, even though the twisted product is not for the Iwahori group.
@@
$\alpha_{\bar{\nu}_{1},\bar{\nu}_{2}}:\mathrm{Grad}_{\bar{\nu}_{1}}(\mathcal{F})\otimes_{\Lambda}^{L}\mathrm{Grad}_{\bar{\nu}_{2}}({\mathcal{G}})\rightarrow\mathrm{Grad}_{\bar{\nu}}({\mathcal{F}}*{\mathcal{G}}).$ (3.53)
@@
$\mathrm{Grad}({\mathcal{F}})\otimes_{\Lambda}^{L}\mathrm{Grad}({\mathcal{G}})\cong\mathrm{Grad}({\mathcal{F}}*{\mathcal{G}})$ (3.54)
@@ 4.1. Background
Here, the first arrow comes from the geometric Satake equivalence of [FS21, Chapter VI], with the Satake category consisting of ula perverse sheaves on ${\operatorname{Hk}}_{G,C}$. The second arrow are just nearby cycles which respect the ula property by Proposition 4.2. Often below, we will find it convenient to still abusively denote by ${\mathscr{Z}}$ the nearby cycles functor $R\Psi\colon{\mathscr{D}}_{\mathrm{ula}}({\operatorname{Hk}}_{G,C})\to{\mathscr{D}}_{\mathrm{ula}}({\operatorname{Hk}}_{{\mathcal{G}},\bar{k}})$.
@@ 4.2. Convolution and fusion
One often finds the expression ${\operatorname{Hk}}_{\mathcal{G}}^{J}={\operatorname{Hk}}_{\mathcal{G}}^{j_{1}}\tilde{\times}\dots\tilde{\times}{\operatorname{Hk}}_{\mathcal{G}}^{j_{n}}$ to denote the convolution Hecke stack. We have already seen that there is a natural correspondence with $n=2$ inducing the convolution product $\ast:{\mathscr{D}}_{\mathrm{\acute{e}t}}({\operatorname{Hk}}_{{\mathcal{G}},S})\times{\mathscr{D}}_{\mathrm{\acute{e}t}}({\operatorname{Hk}}_{{\mathcal{G}},S})\to{\mathscr{D}}_{\mathrm{\acute{e}t}}({\operatorname{Hk}}_{{\mathcal{G}},S})$ for any $S\to\operatorname{{Spd}}O$, see 3.1. We wish to enhance this operation to a monoidal structure of $\infty$-categories in the sense of [Lur17, Definition 4.1.1.10]. This will be quite technical, and we recommend the unaccostumed reader to try to ignore the heavy language at first, and focus on the geometry at hand. After each categorical proof, we also provide an explanation of our constructions at the level of $1$-morphisms of correspondences, which should prove helpful.
@@
Let us recall some of the notions from [Lur17, Section 4.1]. First, we have the (nerve of) the 1-category $\mathrm{Comm}^{\otimes}$ (also denoted by $\mathrm{Fin}_{*}$ or ${\mathbb{E}}^{\otimes}_{\infty}$ in [Lur17]) whose objects are finite pointed sets $\langle n\rangle=\{0,1,\dots,n\}$ with base point 00 and whose morphisms $\langle n\rangle\to\langle m\rangle$ preserve 00. A symmetric monoidal $\infty$-category ${\mathscr{C}}$ is given by a cocartesian fibration ${\mathscr{C}}^{\otimes}\to\mathrm{Comm}^{\otimes}$ of $\infty$-operads, see [Lur17, Example 2.1.2.18], [Lur17, Definition 2.1.1.10], in particular, ${\mathscr{C}}^{\otimes}_{[n]}\simeq{\mathscr{C}}^{n}$ in a natural manner. Similarly, we have the $\infty$-operad $\mathrm{Assoc}^{\otimes}$ (which is equivalent to some other common $\infty$-operads denoted by ${\mathbb{A}}_{\infty}^{\otimes}$ or ${\mathbb{E}}_{1}^{\otimes}$ in [Lur17]) given as the 1-category whose objects are pointed finite set $\langle n\rangle$ and morphisms $\langle n\rangle\to\langle m\rangle$ are pointed maps equipped with a total order on the non-pointed fibers and composition is given by the lexicographical order, see [Lur17, Remark 4.1.1.4]. A monoidal $\infty$-category is a cocartesian fibration ${\mathscr{C}}^{\otimes}\to\mathrm{Assoc}^{\otimes}$ of $\infty$-operads and its underlying $\infty$-category ${\mathscr{C}}$ is the fiber of the fibration over $\langle 1\rangle$. When the cocartesian fibration is clear from the context, we will often abuse language and refer to ${\mathscr{C}}$ as a monoidal $\infty$-category. Such a datum induces by [Lur17, Propositions 2.4.1.7, 2.4.2.5] a map $\mathrm{Assoc}^{\otimes}\to\mathrm{Cat}_{\infty}^{\times}$ that preserves inert morphisms in the sense of [Lur17, Definitions 2.1.1.8, 2.1.2.3] and also cocartesian morphisms, in particular ${\mathscr{C}}^{\otimes}\to\mathrm{Assoc}^{\otimes}$ induces an associative algebra in $\mathrm{Cat}_{\infty}$. Here, $\mathrm{Cat}^{\times}_{\infty}$ denotes $\mathrm{Cat}_{\infty}$ with its cartesian symmetric monoidal structure, cf. [Lur17, Section 2.4.1]. Note that in general maps of $\infty$-operads are not necessarily monoidal, but rather only lax-monoidal, see [Lur17, Definition 2.1.3.7].
@@
One can therefore realize ${\operatorname{Hk}}_{{\mathcal{G}},S}$ as the internal endomorphism object of $[*/L_{S}^{+}{\mathcal{G}}]$ in the $\infty$-category $\mathrm{Corr}(\mathrm{vSt}_{/[*/L_{S}{\mathcal{G}}]})$, and hence it inherits a natural $\infty$-monoidal structure in the category of correspondences by forgetting the slice over $[*/L_{S}{\mathcal{G}}]$. In fact, the functor $\mathrm{Corr}(\mathrm{vSt}_{/[*/L_{S}{\mathcal{I}}]})\to\mathrm{Corr}(\mathrm{vSt})$ is naturally symmetric monoidal. Note that the associated planar $\infty$-operad in the sense of [Lur17, Definition 4.1.3.2] is nothing other than the Čech nerve of the natural map $[*/L_{S}^{+}{\mathcal{G}}]\to[*/L_{S}{\mathcal{G}}]$.
@@
The vertical maps are clearly monoidal and the upper one is fully faithful by [FS21, Proposition VI.9.3], so it suffices to see that the upper diagonal map is monoidal. Indeed, the loop groups $L_{C^{2}}^{+}G$ and $L_{C^{2}}G$ naturally factor into a product away from the diagonal, so we get an induced map ${\mathscr{H}}_{C}^{\otimes}\times{\mathscr{H}}_{C}^{\otimes}\leftarrow{\mathscr{H}}_{C^{2}}^{\otimes,\neq}$ of functors $\mathrm{Assoc}^{\otimes}\to\mathrm{Corr}^{\otimes}(\mathrm{vSt})$ regarded as a correspondence to the left by functoriality of endomorphisms objects. This yields our desired monoidal map upon applying ${\mathscr{D}}^{\otimes}$ and restricting to the monoidal subcategories of perverse sheaves. ∎
@@ 4.3. Associative center
For reasons that will become clear during our induction argument, we replace the exponent $n$ by a finite set $J$ during our proof. If $\lvert J\rvert=1$, this follows already from [GL22, Theorem 4.7] applied to the kimberlite $*_{O_{C}}$, since its reduction equals $*_{k}$ and hence nearby cycles are per definition algebraic, so they can be calculated via the étale site for kimberlites, which is trivial.
If $\lvert J\rvert=2$, then we first compute the stalk of $Rj_{*}{\Lambda}$ at $*_{k\times C}$. We know that partially compactly supported cohomology vanishes by [FS21, Theorem IV.5.3], so $R\Gamma(*_{O_{C}\times C},j_{!}{\Lambda})=0$, compare with [FS21, Proposition V.4.2, Remark V.4.3]. This means that our sought stalk is given by $R\Gamma(*_{C^{2}},{\Lambda})$ which coincides with ${\Lambda}$ thanks to [Sch17, Theorem 19.5]. It remains to compute the stalk at the reduction $*_{k}$ of the kimberlite $*_{O_{C}^{2}}$, so we apply [GL22, Theorem 4.7] once again.
@@
and $K\subset K^{\prime}$ has singleton complement. This implies the claim regarding the stratum $*_{C^{K}}$ again thanks to [Sch17, Theorem 19.5]. As for the non-analytic point $*_{k}$ of $*_{O_{C}^{J}}$, we invoke [GL22, Theorem 4.7] again for the last time. ∎
@@ 4.4. Perversity
Next, we deduce a few important consequences from this theorem. We start by proving that ${\mathscr{Z}}(V)$ is perverse for general parahorics. This is based on a suggestion of Achar to Cass–van den Hoven–Scholbach, see [CvdHS24, Theorem 5.30].
@@
$\mathrm{CT}_{B^{-}}({\mathcal{A}}*{\mathcal{B}})_{\nu}$ $\cong R\Gamma_{c}(S^{-}_{\nu},{\mathcal{A}}*{\mathcal{B}})$ (4.26)
$\cong\bigoplus_{\nu_{1}+\nu_{2}=\nu}R\Gamma_{c}(S^{-}_{\nu_{1}}\tilde{\times}S^{-}_{\nu_{2}},{\mathcal{A}}\tilde{\boxtimes}{\mathcal{B}})$ (4.27)
$\cong\bigoplus_{\nu_{1}+\nu_{2}=\nu}R\Gamma_{c}(S^{-}_{\nu_{1}},{\mathcal{A}})\otimes R\Gamma_{c}(S^{-}_{\nu_{2}},{\mathcal{B}}))$ (4.28)
$\cong\bigoplus_{\nu_{1}+\nu_{2}=\nu}\mathrm{CT}_{B^{-}}({\mathcal{A}})_{\nu_{1}}\otimes\mathrm{CT}_{B^{-}}({\mathcal{B}})_{\nu_{2}}.$ (4.29)
by combining [BR18, Theorem 5.9], [Yu22, Lemma 6.1], and [ALRR24, Corollary 4.16]. This monoidal structure coincides with the natural one on $V\mapsto V_{\hat{T}^{I}}$ under geometric Satake. However, strictly speaking, this construction is not quite complete, because on the one hand [Zhu17, Yu22] works with the Witt vector affine Grassmannian instead of the $B_{\mathrm{dR}}^{+}$-affine Grassmannian, and moreover it would not be immediate that this monoidal structure is compatible with that of [FS21]. In order to fill this gap, we must use the equivalence between the Satake categories over $C$ and $\overline{k}$ for the split group $G_{C}$ in [FS21, VI.6.7] and a theorem of Bando [Ban22, Ban23] showing that it is monoidal, i.e., that pulling back the Satake category of [Zhu17, Yu22] to the context of [FS21] yields the same monoidal structure on perverse sheaves.

### v4

@@ 3.3. Cohomology of Wakimoto filtered objects
By definition ${\mathcal{F}}\ast{\mathscr{I}}_{{\bar{\nu}}}({\Lambda})\cong R{\operatorname{pr}}_{2,\ast}({\operatorname{pr}}_{1}^{\ast}({\mathcal{F}})\otimes^{L}_{\Lambda}\pi^{\ast}\nabla_{t_{\bar{\nu}}})$ and thus we get
@@
This result is analogous to [AR, Proposition 4.5.4]. The existence of a canonical isomorphism follows from Corollary 3.21 by using a filtration of $\mathcal{F}$ by Wakimoto sheaves. Note that the associated graded of ${\mathcal{F}}$ is perverse, so we conclude that, whenever $\bar{\nu}_{1}\prec\bar{\nu}_{2}$, then the cohomology complexes $R\Gamma({\mathrm{Fl}}_{\mathcal{I}},\mathrm{gr}_{\bar{\nu}_{i}}({\mathcal{F}}))$ sit in different degrees with the same parity. This implies that the connecting homomorphisms of the associated long exact sequences vanish. ∎
@@
By Braden’s theorem, the left side of (3.50) naturally identifies with cohomology supported at the corresponding $LU$-orbit ${\mathcal{S}}_{t_{\bar{\nu}}w}$. Since ${\mathcal{S}}_{t_{\bar{\nu}}}$ contains ${\mathrm{Fl}}_{{\mathcal{I}},t_{\bar{\nu}}}$ by Lemma 2.12, the pullback of ${\mathcal{S}}_{t_{\bar{\nu}}w}$ to ${\mathrm{Fl}}_{{\mathcal{I}},\leq t_{\bar{\nu}}}\tilde{\times}{\mathrm{Fl}}_{\mathcal{I}}$ identifies with $L^{+}{\mathcal{U}}\bar{\nu}(\pi)L^{+}{\mathcal{U}}\times^{L^{+}{\mathcal{U}}}{\mathcal{S}}_{w}$. By abuse of notation, we denote the latter scheme by ${\mathrm{Fl}}_{t_{\bar{\nu}}}\tilde{\times}{\mathcal{S}}_{w}$, even though the twisted product is not for the Iwahori group.
@@
$\alpha_{\bar{\nu}_{1},\bar{\nu}_{2}}:\mathrm{Grad}_{\bar{\nu}_{1}}(\mathcal{F})\otimes_{\Lambda}\mathrm{Grad}_{\bar{\nu}_{2}}({\mathcal{G}})\rightarrow\mathrm{Grad}_{\bar{\nu}}({\mathcal{F}}*{\mathcal{G}}).$ (3.53)
@@
$\mathrm{Grad}({\mathcal{F}})\otimes_{\Lambda}\mathrm{Grad}({\mathcal{G}})\cong\mathrm{Grad}({\mathcal{F}}*{\mathcal{G}})$ (3.54)
@@ 4.1. Background
Here, the first arrow comes from the geometric Satake equivalence of [FS21, Chapter VI], with the Satake category consisting of ula perverse sheaves on ${\operatorname{Hk}}_{G,C}$. The second arrow is just nearby cycles which respect the ula property by Proposition 4.2. Often below, we will find it convenient to still abusively denote by ${\mathscr{Z}}$ the nearby cycles functor $R\Psi\colon{\mathscr{D}}_{\mathrm{ula}}({\operatorname{Hk}}_{G,C})\to{\mathscr{D}}_{\mathrm{ula}}({\operatorname{Hk}}_{{\mathcal{G}},\bar{k}})$.
@@ 4.2. Convolution and fusion
One often finds the expression ${\operatorname{Hk}}_{\mathcal{G}}^{J}={\operatorname{Hk}}_{\mathcal{G}}^{j_{1}}\tilde{\times}\dots\tilde{\times}{\operatorname{Hk}}_{\mathcal{G}}^{j_{n}}$ to denote the convolution Hecke stack. We have already seen that there is a natural correspondence with $n=2$ inducing the convolution product $\ast:{\mathscr{D}}_{\mathrm{\acute{e}t}}({\operatorname{Hk}}_{{\mathcal{G}},S})\times{\mathscr{D}}_{\mathrm{\acute{e}t}}({\operatorname{Hk}}_{{\mathcal{G}},S})\to{\mathscr{D}}_{\mathrm{\acute{e}t}}({\operatorname{Hk}}_{{\mathcal{G}},S})$ for any $S\to\operatorname{{Spd}}O$, see 3.1. We wish to enhance this operation to a monoidal structure of $\infty$-categories in the sense of [Lur17, Definition 4.1.1.10]. This will be quite technical, and we recommend the unaccustomed reader to try to ignore the heavy language at first, and focus on the geometry at hand. After each categorical proof, we also provide an explanation of our constructions at the level of $1$-morphisms of correspondences, which should prove helpful.
@@
Let us recall some of the notions from [Lur17, Section 4.1]. First, we have the (nerve of) the 1-category $\mathrm{Comm}^{\otimes}$ (also denoted by $\mathrm{Fin}_{*}$ or ${\mathbb{E}}^{\otimes}_{\infty}$ in [Lur17]) whose objects are finite pointed sets $\langle n\rangle=\{0,1,\dots,n\}$ with base point 00 and whose morphisms $\langle n\rangle\to\langle m\rangle$ preserve 00. A symmetric monoidal $\infty$-category ${\mathscr{C}}$ is given by a cocartesian fibration ${\mathscr{C}}^{\otimes}\to\mathrm{Comm}^{\otimes}$ of $\infty$-operads, see [Lur17, Example 2.1.2.18], [Lur17, Definition 2.1.1.10], in particular, ${\mathscr{C}}^{\otimes}_{[n]}\simeq{\mathscr{C}}^{n}$ in a natural manner. Similarly, we have the $\infty$-operad $\mathrm{Assoc}^{\otimes}$ (which is equivalent to some other common $\infty$-operads denoted by ${\mathbb{A}}_{\infty}^{\otimes}$ or ${\mathbb{E}}_{1}^{\otimes}$ in [Lur17]) given as the 1-category whose objects are pointed finite sets $\langle n\rangle$ and morphisms $\langle n\rangle\to\langle m\rangle$ are pointed maps equipped with a total order on the non-pointed fibers and composition is given by the lexicographical order, see [Lur17, Remark 4.1.1.4]. A monoidal $\infty$-category is a cocartesian fibration ${\mathscr{C}}^{\otimes}\to\mathrm{Assoc}^{\otimes}$ of $\infty$-operads and its underlying $\infty$-category ${\mathscr{C}}$ is the fiber of the fibration over $\langle 1\rangle$. When the cocartesian fibration is clear from the context, we will often abuse language and refer to ${\mathscr{C}}$ as a monoidal $\infty$-category. Such a datum induces by [Lur17, Propositions 2.4.1.7, 2.4.2.5] a map $\mathrm{Assoc}^{\otimes}\to\mathrm{Cat}_{\infty}^{\times}$ that preserves inert morphisms in the sense of [Lur17, Definitions 2.1.1.8, 2.1.2.3] and also cocartesian morphisms, in particular ${\mathscr{C}}^{\otimes}\to\mathrm{Assoc}^{\otimes}$ induces an associative algebra in $\mathrm{Cat}_{\infty}$. Here, $\mathrm{Cat}^{\times}_{\infty}$ denotes $\mathrm{Cat}_{\infty}$ with its cartesian symmetric monoidal structure, cf. [Lur17, Section 2.4.1]. Note that in general maps of $\infty$-operads are not necessarily monoidal, but rather only lax-monoidal, see [Lur17, Definition 2.1.3.7].
@@
One can therefore realize ${\operatorname{Hk}}_{{\mathcal{G}},S}$ as the internal endomorphism object of $[*/L_{S}^{+}{\mathcal{G}}]$ in the $\infty$-category $\mathrm{Corr}(\mathrm{vSt}_{/[*/L_{S}{\mathcal{G}}]})$, and hence it inherits a natural $\infty$-monoidal structure in the category of correspondences by forgetting the slice over $[*/L_{S}{\mathcal{G}}]$. In fact, the functor $\mathrm{Corr}(\mathrm{vSt}_{/[*/L_{S}{\mathcal{G}}]})\to\mathrm{Corr}(\mathrm{vSt})$ is naturally symmetric monoidal. Note that the associated planar $\infty$-operad in the sense of [Lur17, Definition 4.1.3.2] is nothing other than the Čech nerve of the natural map $[*/L_{S}^{+}{\mathcal{G}}]\to[*/L_{S}{\mathcal{G}}]$.
@@
The vertical maps are clearly monoidal and the upper one is fully faithful by [FS21, Proposition VI.9.3], so it suffices to see that the upper diagonal map is monoidal. Indeed, the loop groups $L_{C^{2}}^{+}G$ and $L_{C^{2}}G$ naturally factor into a product away from the diagonal, so we get an induced map ${\mathscr{H}}_{C}^{\otimes}\times{\mathscr{H}}_{C}^{\otimes}\leftarrow{\mathscr{H}}_{C^{2}}^{\otimes,\neq}$ of functors $\mathrm{Assoc}^{\otimes}\to\mathrm{Corr}^{\otimes}(\mathrm{vSt})$ regarded as a correspondence to the left by functoriality of endomorphism objects. This yields our desired monoidal map upon applying ${\mathscr{D}}^{\otimes}$ and restricting to the monoidal subcategories of perverse sheaves. ∎
@@ 4.3. Associative center
For reasons that will become clear during our induction argument, we replace the exponent $n$ by a finite set $J$ during our proof. If $\lvert J\rvert=1$, this follows already from [GL24, Theorem 4.7] applied to the kimberlite $*_{O_{C}}$, since its reduction equals $*_{k}$ and hence nearby cycles are per definition algebraic, so they can be calculated via the étale site for kimberlites, which is trivial.
If $\lvert J\rvert=2$, then we first compute the stalk of $Rj_{*}{\Lambda}$ at $*_{k\times C}$. We know that partially compactly supported cohomology vanishes by [FS21, Theorem IV.5.3], so $R\Gamma(*_{O_{C}\times C},j_{!}{\Lambda})=0$, compare with [FS21, Proposition V.4.2, Remark V.4.3]. This means that our sought stalk is given by $R\Gamma(*_{C^{2}},{\Lambda})$ which coincides with ${\Lambda}$ thanks to [Sch17, Theorem 19.5]. It remains to compute the stalk at the reduction $*_{k}$ of the kimberlite $*_{O_{C}^{2}}$, so we apply [GL24, Theorem 4.7] once again.
@@
and $K\subset K^{\prime}$ has singleton complement. This implies the claim regarding the stratum $*_{C^{K}}$ again thanks to [Sch17, Theorem 19.5]. As for the non-analytic point $*_{k}$ of $*_{O_{C}^{J}}$, we invoke [GL24, Theorem 4.7] again for the last time. ∎
@@
This concludes the part of our work most intensely concerned with $\infty$-categorical matters. It is natural to wonder if these arguments can be equally carried out in equicharacteristic, e.g., in the setting of [AR]. Certainly, it is possible to pass to the corresponding v-sheaves and perform our arguments again, but this seems unsatisfactory. Some care would be needed when replacing $O_{C}^{n}$ by the corresponding scheme-theoretic product, as it does not appear so well-behaved to us. On the other hand, using the nearby cycle construction of [CvdHS24, Definition 3.1] that goes via pushforward along the original curve $\mathbb{A}^{1}_{k}$ and then identifying the right component via $\mathbb{G}_{m,k}$-equivariance, we think our arguments should generalize in a straightforward manner.
@@ 4.4. Perversity
Next, we deduce a few important consequences from this theorem. We start by proving that ${\mathscr{Z}}(V)$ is perverse for general parahorics. This is based on a suggestion of Achar to Cass–van den Hove–Scholbach, see [CvdHS24, Theorem 5.30].
@@
$\mathrm{CT}_{B^{-}}({\mathcal{A}}*{\mathcal{B}})_{\nu}$ $\cong R\Gamma_{c}(S^{-}_{\nu},{\mathcal{A}}*{\mathcal{B}})$
$\cong\bigoplus_{\nu_{1}+\nu_{2}=\nu}R\Gamma_{c}(S^{-}_{\nu_{1}}\tilde{\times}S^{-}_{\nu_{2}},{\mathcal{A}}\tilde{\boxtimes}{\mathcal{B}})$
$\cong\bigoplus_{\nu_{1}+\nu_{2}=\nu}R\Gamma_{c}(S^{-}_{\nu_{1}},{\mathcal{A}})\otimes R\Gamma_{c}(S^{-}_{\nu_{2}},{\mathcal{B}}))$
$\cong\bigoplus_{\nu_{1}+\nu_{2}=\nu}\mathrm{CT}_{B^{-}}({\mathcal{A}})_{\nu_{1}}\otimes\mathrm{CT}_{B^{-}}({\mathcal{B}})_{\nu_{2}}.$
by combining [BR18, Theorem 5.9], [Yu22, Lemma 6.1], and [ALRR24, Corollary 4.16]. This monoidal structure coincides with the natural one on $V\mapsto V_{\hat{T}^{I}}$ under geometric Satake. However, strictly speaking, this construction is not quite complete, because on the one hand [Zhu17, Yu22] works with the Witt vector affine Grassmannian instead of the $B_{\mathrm{dR}}^{+}$-affine Grassmannian, and moreover it would not be immediate that this monoidal structure is compatible with that of [FS21]. In order to fill this gap, we must use the equivalence between the Satake categories over $C$ and $\overline{k}$ for the split group $G_{C}$ in [FS21, VI.6.7] and a theorem of Bando, see [Ban22, Theorem 1.1] showing that it is monoidal, i.e., that pulling back the Satake category of [Zhu17, Yu22] to the context of [FS21] yields the same monoidal structure on perverse sheaves.